OpenStudy (anonymous):

subtract (19x2+5)from(18x2+10)

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

19*2 = 38 and 18*2 = 36 So, (38+5) - (36 + 10) = 43 - 46 = -3

OpenStudy (anonymous):

-x^2-5 it is subtracted FROM...so it means the second one-the 1st one

OpenStudy (anonymous):

Oh damn he's right. My english isn't very good, I'm from Quebec, I didn't understand the question properly. Sorry anthony :(